Q.3 Three year bonds are issued at face value of SR 100, 000 on Jan. 1, 2007, at a stated interest rate of 8%. Calculate the issue price of the bonds assuming a market interest rate of 10%. The present value of SR 1 is .75132 at 10% after 3 years and the present value of an ordinary annuity of SR 1 is 2.48685 at 10% after 3 years.

Prepare discount amortization table and journal entries for three years.
Answer #1
a. Price of bonds is the present value of future cash flows:
Present value of coupons           8,000.00 * 2.48685 = 19,894.80
Present value of face value     1,00,000.00 * 0.75132 = 75,132.00
Present value of cash flows 95,026.80
So, Price of bond is
Working:
Coupon = Par Value * Coupon Rate
=     1,00,000.00 * 8%
=           8,000.00
b. Discount amortization Schedule:
Year Beginning Book Value Interest Expense Coupon Payment Amortization of discount Ending Book Value
a b=a*10% c d=b-c e=a-d
1 95,026.80           9,502.68     8,000.00        1,502.68         96,529.48
2 96,529.48           9,652.95     8,000.00        1,652.95         98,182.43
3 98,182.43           9,818.24     8,000.00        1,818.24     1,00,000.67
Total
c. Journal Entries:
Year Account titles and Explanation Debit Credit
0 Cash 95,026.80
Discount on bonds payable     4,973.20
Bonds Payable 1,00,000.00
(To record issuance of bonds)
1 Interest Expense     9,502.68
Discount on bonds payable        1,502.68
Cash        8,000.00
(To record interest expense)
2 Interest Expense     9,652.95
Discount on bonds payable        1,652.95
Cash        8,000.00
(To record interest expense)
3 Interest Expense     9,818.24
Discount on bonds payable        1,818.24
Cash        8,000.00
(To record interest expense)